Do all of them
for a 22.0-min rest (a) How muth time is spest on the tp Need Help? Ens MacBook Pr control pton command
have 3 days remaining in your WebAasign Grace Perids Tada 52-s time interval starting at the instant ehe brakes Need Helg? ho Assuming conant acceeration, eetermine ow ln the Need Help? t hts the grune MacBook Pro at contro option command
ave 3 days namaining in your WebAusign Grace Porice get.Accsss Tad ta) Find The veliecity of the stject when it is 38.0 m abeve the greand icate the dection with the sign of your ann Lat the ostive crection be upwad lt) Find the total eistance the otjest treveis duing the e Need Help? H Need Helg? MacBook Pr Ls 3t controoption command
Answer #1

3)

Average speed = D/T = 68

D = 68 T ...... (1)

Distance = D = (T - 22/60) x 90.5 ....... (2)

Comparing (1) and (2)

68T = (T - 22/60) x 90.5 = 90.5T - 33.18

T = 1.47 hours

B)

Total distance = (1.47 - 22/60) x 90.5 = 100.27 km

Comment in case any doubt please rate my answer .....

According to the HOMEWORKLIB RULES I can answer one question at a time so please post other questions seperately.....